OpenStudy (cedric):

Which of the following is not a way to represent the solution of the inequality 3(2x - 1) greater than or equal to 4(2x - 3) - 3?

OpenStudy (cedric):

x 6 6 x A number line with a closed circle on 6 and shading to the left. A number line with a closed circle on 6 and shading to the right.

OpenStudy (jiteshmeghwal9):

\[\Huge{3(2x-1)\ge 4(2x-3)-3}\]

OpenStudy (jiteshmeghwal9):

\[\Huge{6x-3 \ge 8x-15}\]

OpenStudy (jiteshmeghwal9):

\[\Huge{6x-8x \ge -15+3}\]

OpenStudy (jiteshmeghwal9):

\[\Huge{-2x \ge -12}\]

OpenStudy (jiteshmeghwal9):

\[\Huge{x \ge 6}\]

OpenStudy (jiteshmeghwal9):

hence , D is the answer :D

OpenStudy (jiteshmeghwal9):

gt it @cedric ?

OpenStudy (cedric):

ok

OpenStudy (cedric):

wait lemme see

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (cedric):

so is it because the 6 times 2 makes the 12?

OpenStudy (jiteshmeghwal9):

yeah!

OpenStudy (cedric):

haha Nice! thanaks man, im giving you a medal

OpenStudy (jiteshmeghwal9):

LOL ! nice to hear :)